1.0 x 102mol of an ideal gas at 3.0 x 102K temperature and 6.0 x 10 6Pa pressure occupies 4.16 x 10-2m3space. Calculate the work done and heat absorbed by the gas if it undergoes expansion isothermally and reversibly to
0.832 m3volume and 3.0 x 10 5Pa pressure.
1
Expert's answer
2019-05-31T00:38:41-0400

Solution.

T = const.


"\Delta U = Q - A"

"\Delta U = 0"

"Q = A"

"A = n \times R \times T \times ln{\frac{V2}{V1}}"

"A = 746836.06 J"

Answer:

A = 747 kJ

Q = 747 kJ